# Singular basins in multiscale systems

### v1

@@ Abstract
Figure 3(a) shows a bistable potential calculated by averaging system (9) with $N=2$ oscillators. Two-dimensional cross-sections of the corresponding singular basin for fixed values of $\varphi_{1}$ and $\mu$ values are illustrated in Figs. 3(b) and (d), respectively. The SF volume vanishes as $\varepsilon$ decreases; figure 3(c) demonstrates this for different values of the detuning parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$. We observe that the scaling of the dependence $V(\varepsilon)$ agrees with the rule (8) for some values of $\Delta_{\omega}$ ($\Delta_{\omega}=0.6$), while for other values, it exhibits a non-monotonic resonant-like behavior ($\Delta_{\omega}=1$ or $\Delta_{\omega}=2.2$). Hence, we conclude that the scaling (8) can fail for the SF in higher-dimensional systems due to the more complex structure of the basin boundaries.
Finally, the singular basin for the system of $N=10$ adaptively coupled rotators (9) is demonstrated in Fig. 4. The figure shows how the SF volume vanishes as a function of $\varepsilon$. Additionally, two trajectories are shown in Fig. 4(b): one starts from the SF and enters rotating motion (red), and one starts very close by that converges to a stable equilibrium.
Figure 3: Illustration of the singular basin in system (9) of $N=2$ coupled phase rotators. (a) The bistable potential of the averaged system. (b) A two-dimensional cross-section of the basin of the full slow-fast system. The notations are the same as in Fig. 2. The cross-section for fixed $\varphi_{1}=1.2461$. (c) The volume of the singular funnel for $\gamma_{c}$, limited to $\mu\in[-10,\,0]$, as a function of $\varepsilon$ for different values of the parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$. (d) The cross-section of the basin for fixed $\mu=2.86$.
Figure 4: Evidence of the singular basin in the system (9) of ten coupled oscillators. (a) Scaling of the singular basin volume $\omega_{i}=-4+(i-1)/9$, $i=1,\dots,10$. Parameters: $\eta=10,\,\alpha=\pi/2,\kappa=1$. (b) Examples of the trajectories: The red trajectory starts from the SF and approaches a rotating attractor asymptotically. The blue trajectory converges to an equilibrium. Initial conditions are $\varphi_{i}(0)=6$, $i=1,\dots,10$, $\mu(0)=-5$ for the red trajectory, and $\mu(0)=-5.1$ for the blue trajectory.
In summary, we described the phenomenon of a singular basin and a singular funnel, which is characteristic of a wide range of slow-fast systems. The presence of a singular basin hinders a reliable timescale reduction of multiscale systems, even when the timescales are significantly different. More specifically, while systems reduced by adiabatic elimination or averaging can exhibit a coexistence of different states with clearly separated basins of attraction, the fast timescale of the original system can lead to the dynamics that is not predicted by the reduced system. This unexpected dynamics looks like a tunnelling transition from one attractor to another when only the reduced system is considered. To understand this phenomenon in greater detail, we developed minimal normal form systems. Additionally, we have demonstrated that singular basins can be observed in higher-dimensional mean-field coupled rotators. While our study focused on bistability, future research involving more than two coexisting attractors and possible tunneling between them would be of interest.
Intriguingly, while any model of a real-world system is necessarily an approximation involving the omission or reduction of faster timescales, our results offer a new perspective on the unexpected transitions of these systems, such as tunnelling between two stable states, which could be triggered by the omitted timescales.
@@ I.2 Numerical averaging in Figure 3
  • We consider the (fast) layer problem of system (5)–(6)], which is given as:
@@ II Supercritical pitchfork normal form with adaptive parameter
$\dot{\mu}=\varepsilon\left(-\mu+aH(\mu)\sqrt{\mu}-b\right),$ (21)
@@ IV.1 Timescale reduction
The slow-fast dynamics of this system for $\alpha=0$ was studied in [Franovic2020], for the deterministic and stochastic case. Here we extend these results (for the deterministic case) to $\alpha\neq 0$.
@@ IV.3 Analytical considerations for the singular funnel scaling
In addition, according to [KrupaSzmolyanExtending], there is only an algebraic contraction of the singular funnel stripe across the fold point. Hence, the exponential estimate (36) remains for the further motion along the fast flow.

### v2

@@ Abstract
Figure 3 shows singular basin and its scaling in system of $N=2$ rotators. The system features the coexistence of two attractors: one stable equilibrium and one periodic or quasi-periodic rotation. A bistable potential calculated by averaging system (9) with $N=2$ oscillators is shown in Figure 3(a). Two-dimensional cross-sections of the corresponding singular basin for fixed values of $\varphi_{1}$ and $\mu$ values are illustrated in Figs. 3(b) and (d), respectively. Figure 3(c) shows how the SF volume shrinks with $\varepsilon$ for different values of the detuning parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$. We observe that $V(\varepsilon)$ follows the scaling (8) for some values of $\Delta_{\omega}$ (e.g. $\Delta_{\omega}=0.6$), while for other values of $\Delta_{\omega}$, it exhibits resonance-like deviations from (8) (e.g. $\Delta_{\omega}=1$ or $\Delta_{\omega}=2.2$). Hence, we conclude that higher-dimensional systems can have SFs that deviate from the simple scaling (8) and leave this as a question for future research.
Figure 4: (a) The volume of the basin of attraction of ten adaptive phase oscillators (9-10) for $N=10$, limited to $10\leq\mu\leq 0$, as a function of $\varepsilon$ obtained using Monte Carlo simulation see the supplemental material fro more details. Its decrease with $\varepsilon$ indicates the presence of a singular funnel (SF). (b,c) Two trajectories starting from close initial conditions: $\varphi_{i}(0)=6$ for $i=1,\dots,10$, with $\mu(0)=-5$ (red) and $\mu(0)=-5.1$ (blue). The red trajectory starts from SF and therefore it is attracted to a stable rotating state. Other parameters: $\omega_{i}=-4+(i-1)/9$, $i=1,\dots,10$, $\eta=10$, $\alpha=\pi/2$, and $\kappa=1$.
A visualization of SFs in high-dimensional systems via two-dimensional cross sections is very difficult. Instead, the estimated volume of the SF for the case of $10$ coupled rotators (eleven-dimensional phase space) is plotted in Figure 4(a). Similar to the lower-dimensional examples, the volume shrinks with $\varepsilon$, indicating the presence of a SF. Additionally, Fig. 4(b) shows two trajectories: (red) one starting from the SF and converging to periodic rotations, and (blue) one starting and staying very close by, but eventually converging to the stable equilibrium $e_{1}$, which is another indication of a SF.
Finally, singular basins are a robust phenomenon. This is due to the fact that its ingredients persist under small parameter changes. This applies to the stable manifolds of the saddle equilibria and the coexisting attractors. We illustrate the parameter region of the singular basin and the possible mechanisms of its destruction for system (56) in the supplemental material.
In summary, we have uncovered singular basins with singular funnels (SFs) in bistable slow-fast systems, and showed that they are robust and ubiquitous. The SFs prevent useful and widely used dimensionality reductions, even when the timescales differ significantly and there is a unique stable quasistatic state that changes continuously with the slow variable(s). Although adiabatic elimination or averaging of the fast variables retains the coexistence of different stable states and different basins of attraction in the reduced slow subsystem, the fast dynamics of the full system can lead to transitions between different stable states via SFs that are not possible in the slow subsystem. From the slow subsystem’s perspective, these unexpected dynamics resemble quantum-like ‘tunnelling’ or ‘teleporting’ from one attractor to another. We have constructed low-dimensional canonical systems that can be considered normal forms for a singular basin and showed a universal scaling of SFs with increasing timescale ratio in these systems. Additionally, we have demonstrated that SFs can form intricate structures in the phase space of higher-dimensional systems with slowly adaptive mean-field coupling. A challenge for future research is to understand singular basins in systems with more than two timescales and more than two coexisting stable states.
Perhaps most intriguingly of all, any mathematical model of a real-world system is necessarily an approximation that omits some fast variables. Thus, our results also provide new insight into the potential for unexpected critical transitions [3]. For example, transitions between different stable states, which are not captured by the models due to omitting some fast variables, could be triggered in real systems.
@@ I.2 Numerical averaging in Figure 3
  • We consider the (fast) layer problem of system (5)–(6), which is given as:
@@ II Supercritical pitchfork normal form with adaptive parameter
$\dot{\mu}=\varepsilon\left(-\mu+a\sqrt{\mu H(\mu)}-b\right),$ (21)
@@ IV.1 Timescale reduction
The slow-fast dynamics of this system for $\alpha=0$ was studied in [18], for the deterministic and stochastic case. Here we extend these results (for the deterministic case) to $\alpha\neq 0$.
@@ IV.3 Analytical considerations for the singular funnel scaling
In addition, according to [31], there is only an algebraic contraction of the singular funnel stripe across the fold point. Hence, the exponential estimate (36) remains for the further motion along the fast flow.